# Exploration and Incentivizing Participation in Clinical Trials

### v6

@@ 2 Our model: incentivized clinical trial
We consider a clinical trial (mechanism) with some patients (agents). The mechanism interacts with the agents as a multi-armed bandit algorithm, but optimizes a very different statistical objective, and operates under economic constraints to ensure that the agents are incentivized to participate. Accordingly, we split the model description into three parts: the bandit model for the interaction protocol of the mechanism (Section 2.1), the statistical model for its performance objective (Section 2.2), and the economic model for the agents’ incentives (Section 2.3).
We first present the model for homogeneous agents (i.e., agents with the same “type”), and then extend it to heterogeneous agents in Section 6. This is to disentangle the core model of incentivized clinical trial, which is quite complex even for homogeneous agents, from the subtlety and heavy notation associated with multiple agent types.
@@ 2.1 Bandit model: the interaction protocol
    a new agent arrives, with outcomes $\left(\,\omega_{a,t}\in\Omega:\,a\in[n]\,\right)$ that are unknown to the mechanism.
@@
    The mechanism chooses the sampling distribution over arms, denoted by $p_{t}$, samples an arm $a=a_{t}$ independently at random from this distribution, and assigns this arm to the agent.
@@
    The outcome $\omega_{t}=\omega_{a,t}\in\Omega$ is realized and observed by the mechanism and the agent.
Each outcome $\omega$ is assigned a score $f(\omega)\in[0,1]$, which may e.g., represent this outcome’s value for the society. The scoring function $f:\Omega\to[0,1]$ is exogenously fixed and known. Note that the mechanism follows a standard protocol of multi-armed bandits, in the sense that in each round it chooses an arm $a_{t}$ and observes a numerical score $f(\omega_{t})$ for playing this arm.
The outcome table $\left(\,\omega_{a,t}\in\Omega:\,a\in[n]\,\right)_{t\in[T]}$ is fixed in advance (where round $t$ indexes rows). We posit that this entire table is drawn from some distribution, which we henceforth call an adversary. [^6] This corresponds to “oblivious adversary” in multi-armed bandits, as if there is an adversary who chooses the outcome table upfront. An analogous game-theoretic terminology is a non-strategic player called “Nature”. An important special case is stochastic adversary, whereby each row $t$ of the table is drawn independently from the same fixed distribution.
@@
We posit that the agents cannot change the arms assigned by the mechanism. Indeed, a patient who voluntarily participates in a clinical trial has no reason to change her assigned treatment simply because she cannot observe what this treatment is.
@@ 2.2 Statistical model: the performance objective
More precisely, the mechanism designates the first $T_{0}$ rounds as a warm-up stage, for some $T_{0}\leq T/2$ chosen in advance. We are interested in the average score over the remaining rounds, which constitute the main stage. (The warm-up stage corresponds to a smaller-scale pilot experiment; we use the data from this stage to create incentives for the main stage.)
@@
The meaning of the scoring function $f:\Omega\to[0,1]$ is as follows. First, $f_{\mathtt{adv}}(a)$ is the frequency of a given outcome $\omega^{*}\in\Omega$ in the main stage if the scoring function is given by $f(\omega)={\bf 1}_{\left\{\,\omega=\omega^{*}\,\right\}}$. Alternatively, if $f(\omega)$ is the outcome’s value for the society, then $f_{\mathtt{adv}}(a)$ is the average value of arm $a$. We keep $f$ abstract throughout, and handle all scoring functions in a uniform way. In particular, the mechanism’s choice of sampling distributions $p_{t}$ does not depend on $f$.
After the last round, the mechanism outputs an estimate $\widehat{f}_{\mathtt{mech}}(a)\in[0,1]$ for $f_{\mathtt{adv}}(a)$, for each arm $a$. The estimation error is defined as the largest mean squared error (MSE). Formally, for mechanism $\mathtt{mech}$ and adversary $\mathtt{adv}$ we define
@@
The full specification of our mechanism is that it chooses $T_{0}$, then follows the interaction protocol from Section 2.1, and finally computes the estimates $\widehat{f}_{\mathtt{mech}}(a)$, $a\in[n]$. Thus, the performance objective is to minimize (2.2) uniformly over all adversaries $\mathtt{adv}$.
@@
What makes our problem non-trivial are the economic constraints. Without them, a trivial solution for the stated performance objective is to take $T_{0}=0$ and randomize uniformly over arms, independently in each round. Indeed, this is a solution commonly used in practice. Then one can use a standard IPS estimator and achieve $\mathtt{ERR}(\mathtt{mech}\mid\mathtt{adv})\leq n/T$ (see Eqns. (3.2) and (3.3)).
@@ 2.3 Economic model: agents, beliefs and incentives
The agents’ information structure is as follows. (i) Each agent $t$ does not observe its own outcome tuple $\left(\,\omega_{a,t}:\,a\in[n]\,\right)$. (ii) Agent $t$ does not observe anything about the previous rounds. This is a reasonable assumption in clinical trials: indeed, not revealing any information about an ongoing trial is strongly encouraged (e.g.,  Arango et al.,, 2012). (iii) Agent $t$ does not observe the chosen distribution $p_{t}$ or arm $a_{t}$. Indeed, clinical trials typically present different treatments in an identical way, e.g., as a pill that looks and feels exactly the same; this is strongly encouraged whenever possible (e.g.,  Arango et al.,, 2012). (iv) Each agent knows the specification of the mechanism before (s)he arrives. In practical terms, the entity which runs the clinical trial (e.g., a hospital) commits to a particular design of the trial, and this design is revealed to all patients. This property typically holds in practice, as it is enforced via regulation.
@@
While (2.3) is a natural “outside option” within our model, our analysis accommodates an arbitrary, exogenously given $U_{\mathtt{out}}\in[0,1]$, and all results carry over with minimal modifications. This allows us to model several effects. First, some medical treatments might not be available outside of a clinical trial. Then we could take the $\max$ in (2.3) only over the treatments that are. Conversely, a standard treatment might not be available inside the trial, e.g., in a new treatment vs. placebo trials. Then the $\max$ in (2.3) should also include the standard treatment. Finally, patients might assign positive utility to participation itself, e.g., because they value contributing to science or they expect to receive higher-quality medical care. This can be modeled (albeit in a rather idealized way) by decreasing $U_{\mathtt{out}}$ in (2.3) by some known amount.
3 Preliminaries
@@
Non-adaptivity. Our results focus on mechanisms that do not adapt to new observations over some time period. Given a contiguous set $S$ of rounds, we say that a mechanism is non-data-adaptive over $S$ if for each round $t\in S$ the sampling distribution $p_{t}$ does not depend on the data collected during $S$. If furthermore the mechanism uses the same recommendation policy $\sigma_{t}$ throughout $S$, then we call it stationary over $S$. In this terminology, all mechanisms that we design are stationary over the entire main stage. Our lower bounds merely posit that the mechanism is not data-adaptive on some $S$ such that $|S|/T$ is an absolute constant.
@@
Consider a mechanism $\mathtt{mech}$ which chooses a fixed sequence of sampling distributions $\left(\,p_{t}:t\in S\,\right)$. If $\mathtt{mech}$ uses the IPS estimator, $\widehat{f}_{\mathtt{mech}}=\mathtt{IPS}$, then [^10] The guarantee (3.3) is folklore. It is essentially optimal for any estimator $\widehat{f}_{\mathtt{mech}}$. See Appendix C for proofs.
@@ 4 Homogeneous agents
This section treats the paradigmatic case of homogeneous agents We obtain an optimal solution, with nearly matching upper and lower bounds, and illustrate the key techniques.
Benchmark. We express our results in terms of a benchmark which encapsulates the optimal dependence on the prior. To this end, we consider recommendation policies that input the true state ($\mathtt{sig}=\psi^{*}$); we call them state-aware. We optimize among all BIR, state-aware policies $\sigma$ so as to minimize the error upper bound (3.3). More precisely, we’d like to minimize $\min_{a}1/\sigma_{a}(\psi^{*})$ in the worst case over adversaries $\mathtt{adv}$, i.e., in the worst case over all possible realizations of $\psi^{*}$.
@@
If $U^{*}_{\mathcal{P}}>U_{\mathtt{out}}$ then $\mathtt{bench}(\mathcal{P})<\infty$.  [^12] Condition $U^{*}_{\mathcal{P}}>U_{\mathtt{out}}$ states that $a^{*}(\psi^{*})$ is strictly BIR as a state-aware recommendation policy.
Our mechanism. For the main stage, we use a recommendation policy similar to epsilon-greedy, a standard (but suboptimal) bandit algorithm. Namely, fixing $\epsilon\geq 0$ and state $\psi$, $(\epsilon,\psi)\text{-$\mathtt{greedy}$}$ is a recommendation policy that explores with probability $\epsilon\geq 0$, choosing an arm independently and uniformly at random, and exploits with the remaining probability, choosing the best arm $a^{*}(\psi)$ for a given state $\psi$. Our mechanism uses policy $(\epsilon,\psi)\text{-$\mathtt{greedy}$}$ throughout the main stage, for some fixed $\epsilon,\psi$ defined below. Since the true state $\psi^{*}$ is not known, we instead use state $\psi=\overline{\psi}$ which summarizes the data from the warm-up stage. Specifically, let $\overline{\psi}(a)$ be the empirical distribution over outcomes observed in the rounds of the warm-up stage when a given arm $a$ is chosen. [^13] The average state $\overline{\psi}(a)$ is not necessarily in $\mathtt{support}(\mathcal{P})$, but the $(\epsilon,\overline{\psi})\text{-$\mathtt{greedy}$}$ policy is well-defined.
Our positive result is conditional on collecting enough samples in the warm-up stage.
@@
For analysis, let’s first characterize $\mathtt{bench}(\mathcal{P})$ via the epsilon-greedy policy $(\epsilon,\psi^{*})\text{-$\mathtt{greedy}$}$.
Lemma 4.3.
Proof.
The first inequality in (LABEL:eq:lm:pf:bench-single) holds since $\sigma_{a}(\psi)\geq\epsilon/n$, so for each state the right-hand side shifts probabilities towards arms with weakly smaller expected utility. It follows that policy $\sigma^{\prime}$ is BIR. So, this policy gets within $\delta$ from the supremum in (4.3). This holds for any $\delta>0$. ∎
Proof of Claim 4.1.
@@
(All expectations are over $\psi\sim\mathcal{P}$.) The first inequality holds by (4) (and the fact that utilities are at most $1$). The last inequality holds because $\epsilon\cdot\mathtt{gap}(\mathcal{P})-p_{\mathtt{err}}-2\delta\geq 0$ by our choice of $\epsilon,\delta$.
@@
Finally, recall that in each round of the main stage, each arm is sampled with probability at least $\epsilon/n=1/(2\cdot\mathtt{bench}(\mathcal{P}))$. Thus, Theorem 4.2(a) follows from (3.3). ∎
Warm-up stage. We rely on prior work on incentivized exploration to guarantee warm-up samples. Restated in our notation, the relevant guarantee is as follows:
@@
    if $M_{\mathcal{P}}=\infty$ and $n=2$ then no strictly BIR mechanism can sample both arms.
Thus, sufficient warm-up stage for Theorem 4.2 lasts $T_{0}=M_{\mathcal{P}}\cdot N_{\mathcal{P}}$ rounds, and consists of $N_{\mathcal{P}}$ consecutive runs of the mechanism from Lemma 4.4(a).
@@
Lower bounds. We prove that the guarantee in Theorem 4.2(a) is essentially optimal, in the worst case over the adversaries. Our result applies to all mechanisms that are non-data-adaptive over some constant (but possibly small) fraction of rounds, but are unrestricted otherwise.
@@
$\mathtt{ERR}\left(\,\mathtt{mech}\mid\mathtt{adv}\,\right)=\Omega\left(\,\mathtt{bench}(\mathcal{P})/T\,\right).$ (4.4)
Thus, Theorem 4.5 allows per-round policies that change over time and can themselves be jointly chosen at random. In rounds $t\not\in S$ the policies can depend on all past observations. Importantly, this includes “multi-stage” mechanisms that partition the time into a constant number of contiguous “stages” and are non-data-adaptive within each (possibly with a data-adaptive warm-up stage of duration, say, $<T/2$). Most medical trials in practice fit this model, see Footnote 3.
@@
$\mathtt{ERR}\left(\,\mathtt{mech}\mid\mathtt{adv}^{\dagger}\,\right)\geq\textstyle\Omega\left(\,\frac{1}{|S|^{2}}\;\max_{a\in[n]}\;\sum_{t\in S}\min\left\{\,1/\operatornamewithlimits{\mathbb{E}}\left[\,p_{t}(a)\,\right],\sqrt{|S|}\,\right\}\,\right),$ (4.5)
for some adversary $\mathtt{adv}^{\dagger}$ determined by the expected sampling probabilities $\left(\,\operatornamewithlimits{\mathbb{E}}[p_{t}]:\,t\in S\,\right)$.  [^14] Eq. (4.5) is a general statistical tool, see Appendix C for a standalone formulation, discussion, and proof.
@@
$\mathtt{ERR}\left(\,\mathtt{mech}\mid\mathtt{adv}\,\right)$ $\geq\textstyle\Omega\left(\,\frac{1}{|S|^{2}}\sum_{t\in S}\;\max\left\{\,\tfrac{1}{q_{t}},\;\sqrt{|S|}\,\right\}\,\right),\qquad\text{where }q_{t}:=\operatornamewithlimits{\mathbb{E}}\left[\,p_{t}(a_{0})\,\right]$
$\geq\textstyle\Omega\left(\,\frac{1}{|S|^{2}}\;\left(\,\sqrt{|S|}\cdot|S_{\mathtt{L}}|+\sum_{t\in S_{\mathtt{H}}}\;\tfrac{1}{q_{t}}\,\right)\,\right),$ (4.7)
@@
By harmonic-arithmetic mean inequality, $\sum_{t\in S_{\mathtt{H}}}\tfrac{1}{q_{t}}\geq|S_{\mathtt{H}}|^{2}/N_{\mathtt{H}}\geq|S_{\mathtt{H}}|^{2}\cdot\mathtt{bench}(\mathcal{P})/|S|$. Plugging this back into (4.7) and noting that $|S_{\mathtt{L}}|+|S_{\mathtt{H}}|=|S|\geq T/2$, we have
$\mathtt{ERR}\left(\,\mathtt{mech}\mid\mathtt{adv}\,\right)$ $\geq\textstyle\Omega\left(\,\frac{1}{|S|^{2}}\;\left(\,\sqrt{|S|}\cdot|S_{\mathtt{L}}|+|S_{\mathtt{H}}|^{2}\cdot\mathtt{bench}(\mathcal{P})/|S|\,\right)\,\right)\geq\Omega\left(\,\mathtt{bench}(\mathcal{P})/T\,\right).$
5 Heterogeneous agents: an overview
Agent heterogeneity is two-fold: different agent types may have different outcome distributions for the same arm, and different subjective (and not directly observable) utilities for the same outcome. We model this as follows: public types determine outcome distributions and the private types determine subjective utilities. [^15] So, each agent’s type consists of the public type and the private type; only the former is known to the mechanism. The intuition is that beliefs about objective outcomes are driven by the patient’s objective medical history, which is typically well-documented and available to the clinical trial. However, patients’ preferences over outcome distributions can be highly subjective, e.g., in evaluating the disutility of particular medical complications. These subjective preferences are not known to the mechanism. In our model, each agent reports its private type upon arrival, and the mechanism must incentivize truthful reporting (in addition to the BIR property). [^16] We express this new requirement in a Bayesian framework, and term it Bayesian Incentive-Compatibility, or BIC.
We study this setting in Section 6. We obtain matching upper and lower bounds like in Section 4, but with a new benchmark, mechanism, and analysis. This novelty is necessitated by the BIC condition which binds the types together. [^17] Indeed, without private types one can use mechanism from Section 4 tailored to each particular type. We define a new benchmark, denoted $\mathtt{bench}(\mathcal{P},\Theta)$, which generalizes (4.1) to a given set $\Theta$ of agent types. Compared to the expression in (4.1), this benchmark requires all BIC and BIR per-round policies under the $\inf$, and takes the $\max$ over $\Theta$ (as well as over all states). Our mechanism attains the same error bound as in Theorem 4.2 relative to $\mathtt{bench}(\mathcal{P},\Theta)$, with a similar but much more technical sufficient condition for BIR. The lower bound carries over from Theorem 4.5, replacing $\mathtt{bench}(\mathcal{P})$ with $\mathtt{bench}(\mathcal{P},\Theta)$.

### v8

@@ 2 Our model: incentivized clinical trial
We first present our model for homogeneous agents (i.e., agents with the same “type”), then extend it to heterogeneous agents in Section 4. This is to disentangle the core model, which is quite complex even for a single agent type, from some subtlety and heavy notation needed for multiple types.
Thus, we consider a randomized clinical trial (RCT) with some patients, interpreted as a mechanism) and agents, respectively. The mechanism interacts with the agents as a multi-armed bandit algorithm, but optimizes a very different statistical objective, and operates under economic constraints to incentivize participation. Accordingly, we split the model description into three subsections: the bandit model for the interaction protocol the statistical model for its performance objective, and the economic model for the agents’ incentives.
@@ 2.1 Bandit model: the interaction protocol
    a new agent arrives
@@
    The mechanism chooses the sampling distribution over arms, denoted by $p_{t}$, samples an arm $a_{t}\in[n]$ independently at random from this distribution, and assigns this arm to the agent.
@@
    The outcome $\omega_{t}\in\Omega$ is realized and observed by the mechanism and the agent.
Each outcome $\omega$ is assigned a score $f(\omega)\in[0,1]$, which may e.g., represent this outcome’s value. The scoring function $f:\Omega\to[0,1]$ is exogenously fixed and known (its role is explained in Remark 2.2). Note that the mechanism follows a standard protocol of multi-armed bandits: in each round it chooses an arm $a_{t}$ and observes a numerical score $f(\omega_{t})$ for playing this arm.
The outcome table $\left(\,\omega_{a,t}\in\Omega:\,a\in[n],\;t\in[T]\,\right)$ is fixed in advance (where round $t$ indexes rows), so that $\omega_{t}=\omega_{a_{t},\;t}\in\Omega$ for each round $t$. We posit that this entire table is drawn from some distribution, which we henceforth call an adversary. [^12] This corresponds to “oblivious adversary” in multi-armed bandits, as if there is an adversary who chooses the outcome table upfront. An analogous game-theoretic terminology is a non-strategic player called “Nature”. An important special case is stochastic adversary, whereby each row $t$ of the table is drawn independently from the same fixed distribution.
@@
The agents cannot change the assigned arms. Indeed, Indeed, an RCT participant cannot change her assigned treatment and cannot even observe what this treatment is.
@@ 2.2 Statistical model: the performance objective
More precisely, the mechanism designates the first $T_{0}$ rounds as a warm-up stage, a small-scale pilot experiment to create incentives, for some $T_{0}\leq T/2$ chosen in advance. We are interested in the average score over the remaining rounds, which constitute the main stage.
@@
After round $T$, the mechanism outputs an estimate $\widehat{f}_{\mathtt{mech}}(a)\in[0,1]$ for $f_{\mathtt{adv}}(a)$, for each arm $a$.
Remark 2.2.
We focus on the following two scenarios. First, $f_{\mathtt{adv}}(a)$ is the frequency of a given outcome $\omega^{*}\in\Omega$ in the main stage; then the scoring function is $f(\omega)={\bf 1}_{\left\{\,\omega=\omega^{*}\,\right\}}$. Second, if $f(\omega)$ is the outcome’s value (for the agent and/or the society), then $f_{\mathtt{adv}}(a)$ is the average value of arm $a$ [^13] In particular, the average treatment effect can be expressed as $f_{\mathtt{adv}}(\text{treatment})-f_{\mathtt{adv}}(\text{control})$. We keep $f$ abstract throughout to handle both scenarios in a uniform way.
The estimation error is defined as the largest mean squared error (MSE). Formally, for mechanism $\mathtt{mech}$ and adversary $\mathtt{adv}$ we define
@@
The full specification of our mechanism is that it chooses $T_{0}$, then follows the interaction protocol from Section 2.1, and finally computes the estimates $\widehat{f}_{\mathtt{mech}}(a)$, $a\in[n]$. Thus, the performance objective is to minimize (2.2) uniformly over all adversaries $\mathtt{adv}$. Note that (2.2) essentially requires the mechanism to handle all $f$ at once.
@@
Without economic constraints, a trivial solution is to take $T_{0}=0$ and randomize uniformly over arms, independently in each round. Indeed, this is a solution commonly used in practice. A standard IPS estimator (see Section 2.4) achieves $\mathtt{ERR}(\mathtt{mech}\mid\mathtt{adv})\leq n/T$.
@@ 2.3 Economic model: agents, beliefs and incentives
The agents’ information structure is as follows. (i) Each agent $t$ does not observe anything about the previous rounds. Indeed, not revealing any information about an ongoing trial is strongly encouraged in practice (e.g.,  Arango et al.,, 2012). Consequently, agent $t$ does not observe the chosen distribution $p_{t}$ if it depends on the history. (ii) Each agent $t$ does not observe the assigned arm $a_{t}$. Indeed, RCTs typically present different treatments in an identical way, e.g., as a pill that looks and feels exactly the same; this is strongly encouraged whenever possible (e.g.,  Arango et al.,, 2012). (iii) Each agent knows the specification of the mechanism before (s)he arrives. In practical terms, the entity which runs the RCT (e.g., a hospital) commits to a particular design of the trial, and this design is revealed to all patients. This property is typically enforced via regulation.
@@
While (2.3) is a natural “outside option” within our model, our analysis accommodates an arbitrary, exogenously given $U_{\mathtt{out}}\in[0,1]$, and all results carry over with minimal modifications. This allows us to model several effects. First, some medical treatments might not be available outside of a clinical trial. Then we could take the $\max$ in (2.3) only over the treatments that are. Conversely, a standard treatment might not be available inside the trial, e.g., in a new treatment vs. placebo trials. Then the $\max$ in (2.3) should also include the standard treatment. Finally, patients might assign positive utility to participation itself, e.g., because they value contributing to science or they expect to receive higher-quality and/or less expensive medical care. This can be modeled (albeit in a rather idealized way) by decreasing $U_{\mathtt{out}}$ in (2.3) by some known amount.
2.4 Preliminaries
@@
Consider a mechanism $\mathtt{mech}$ which chooses a fixed sequence of sampling distributions $\left(\,p_{t}:t\in S\,\right)$. If $\mathtt{mech}$ uses the IPS estimator, $\widehat{f}_{\mathtt{mech}}=\mathtt{IPS}$, then
@@ 3 Homogeneous agents
This section treats the fundamental case of homogeneous agents. We obtain an optimal solution, with nearly matching upper/lower bounds, and introduce key ideas for the general case.
3.1 Benchmark
We express our results in terms of a benchmark which encapsulates the optimal dependence on the prior. To this end, we consider recommendation policies that input the true state ($\mathtt{sig}=\psi^{*}$); we call them state-aware. We optimize among all BIR, state-aware policies $\sigma$ so as to minimize the error upper bound (2.6). More precisely, we’d like to minimize $\min_{a}1/\sigma_{a}(\psi^{*})$ in the worst case over adversaries $\mathtt{adv}$, i.e., in the worst case over all possible realizations of $\psi^{*}$.
@@ 3.1 Benchmark
An alternative characterization of this benchmark is based on epsilon-greedy, a well-known bandit algorithm captured (in our terms) by the next definition.
Definition 3.1.
Fixing $\epsilon\geq 0$ and state $\psi$, $(\epsilon,\psi)\text{-$\mathtt{greedy}$}$ is a recommendation policy that explores with probability $\epsilon\geq 0$, choosing an arm independently and uniformly at random, and exploits with the remaining probability, choosing the best arm $a^{*}(\psi)$ for a given state $\psi$.
We use this policy with $\psi=\psi^{*}$ and the largest $\epsilon$ such that this policy is BIR.
Lemma 3.2.
Remark 3.3.
Both formulations, (3.1) and (3.2), are used in the analysis: the former for the lower bound and the latter for the upper bound. $(\epsilon,\psi)\text{-$\mathtt{greedy}$}$ policy is also used in our mechanism. While epsilon-greedy is suboptimal as a regret-minimizing bandit algorithm, it suffices to achieve optimal performance for our purposes.
Proof of Lemma 3.2.
The inequality in (LABEL:eq:lm:pf:bench-single) holds since $\sigma_{a}(\psi)\geq\epsilon/n$, so for each state the right-hand side shifts probabilities towards arms with weakly smaller expected utility. It follows that policy $\sigma^{\prime}$ is BIR. So, this policy gets within $\delta$ from the supremum in (3.2). This holds for any $\delta>0$. ∎
@@
If $U^{*}_{\mathcal{P}}>U_{\mathtt{out}}$ then $\mathtt{bench}(\mathcal{P})<\infty$.
Proof.
3.2 Our mechanism: Main stage
We focus on the main stage, coming back to the warm-up stage in Section 3.3. We use the same recommendation policy in all rounds of the main stage. Namely, we use policy $(\epsilon,\psi)\text{-$\mathtt{greedy}$}$, for some fixed $\epsilon,\psi$. Since the true state $\psi^{*}$ is not known, we instead use state $\psi=\overline{\psi}$ which summarizes the data from the warm-up stage. Specifically, let $\overline{\psi}(a)$ be the empirical distribution over outcomes observed in the rounds of the warm-up stage when a given arm $a$ is chosen. Note that the average state $\overline{\psi}(a)$ is not necessarily in $\mathtt{support}(\mathcal{P})$, but the $(\epsilon,\overline{\psi})\text{-$\mathtt{greedy}$}$ policy is well-defined.
To recap, our mechanism uses $(\epsilon,\overline{\psi})\text{-$\mathtt{greedy}$}$ recommendation policy in all rounds, for some $\epsilon>0$. Our guarantee is conditional on collecting enough samples in the warm-up stage.
@@ 3.2 Our mechanism: Main stage
We illustrate $\mathtt{bench}(\mathcal{P})$ and $N_{\mathcal{P}}$ with a simple numerical example in Appendix A.
@@
where we denote policy $(2\epsilon,\psi^{*})\text{-$\mathtt{greedy}$}$ with $\sigma^{*}$. The first inequality holds by (3.2) (and the fact that utilities are at most $1$). The last inequality holds because $\epsilon\cdot\mathtt{gap}(\mathcal{P})-p_{\mathtt{err}}-2\delta\geq 0$ by our choice of $\epsilon,\delta$.
@@
Finally, recall that in each round of the main stage, each arm is sampled with probability at least $\epsilon/n=1/(2\cdot\mathtt{bench}(\mathcal{P}))$. Thus, Theorem 3.5(a) follows from Eq. (2.6). ∎
3.3 Our mechanism: Warm-up stage
We rely on prior work on incentivized exploration to guarantee warm-up samples. Restated in our notation, the relevant guarantee is as follows:
@@ 3.3 Our mechanism: Warm-up stage
    if $M_{\mathcal{P}}=\infty$ and $n=2$ arms then no strictly BIR mechanism can sample both arms.
Thus, sufficient warm-up stage for Theorem 3.5 lasts $T_{0}=M_{\mathcal{P}}\cdot N_{\mathcal{P}}$ rounds, and consists of $N_{\mathcal{P}}$ consecutive runs of the mechanism from Lemma 3.6(a). The expression for $M_{\mathcal{P}}$ is somewhat tedious, and not essential for this paper. [^18] However, let us present it for $n=2$ arms, for the sake of completeness. Consider $\Delta_{\psi}:=U_{\psi}(1)-U_{\psi}(2)$, the “gap” between the arms’ expected utilities. W.l.o.g., assume that $\operatornamewithlimits{\mathbb{E}}_{\psi}\left[\,\Delta_{\psi}\,\right]\geq 0$, i.e., arm $1$ is (weakly) preferred initially. Let $X_{(k)}:=\operatornamewithlimits{\mathbb{E}}_{\psi}\left[\,-\Delta_{\psi}\mid S_{(k)}\,\right]$, where $S_{(k)}$ is a tuple of $k$ independent reward samples from arm $1$. Then, suppose there exists some finite $k=k_{\mathcal{P}}$ such that $\Pr\left[\,X_{(k)}>0\,\right]>0$ (if not, set $M_{\mathcal{P}}=\infty$). Let $Y=X_{(k_{\mathcal{P}})}$. Then $M_{\mathcal{P}}=(n+1)\max\left(\,k_{\mathcal{P}},\;\frac{\operatornamewithlimits{\mathbb{E}}_{\psi}\left[\,\Delta_{\psi}\,\right]}{\operatornamewithlimits{\mathbb{E}}\left[\,Y\mid Y>0\,\right]\;\Pr[Y>0]}\,\right).$
@@
3.4 Lower bound
We prove that the guarantee in Theorem 3.5(a) is essentially optimal in the worst case.
Our result applies to all mechanisms that do not adapt to new observations over some constant (but possibly small) fraction of rounds, and are unrestricted otherwise. Formally, given a contiguous set $S$ of rounds, we say that a mechanism is non-data-adaptive over $S$ if for each round $t\in S$ the sampling distribution $p_{t}$ does not depend on the data collected during $S$.
@@ 3.4 Lower bound
$\mathtt{ERR}\left(\,\mathtt{mech}\mid\mathtt{adv}\,\right)=\Omega\left(\,\frac{\mathtt{bench}(\mathcal{P})}{T}\,\right).$ (3.4)
Thus, Theorem 3.7 allows per-round policies that change over time and can themselves be jointly chosen at random. In rounds $t\not\in S$ the policies can depend on all past observations. Importantly, this includes “multi-stage” mechanisms that partition the time into a constant number of contiguous “stages” and are non-data-adaptive within each (possibly with a data-adaptive warm-up stage of duration, say, $<T/2$). Most medical trials in practice fit this model, see Footnote 7.
@@
[…]
for some adversary $\mathtt{adv}^{\dagger}$ determined by the expected sampling probabilities $\left(\,\operatornamewithlimits{\mathbb{E}}[p_{t}]:\,t\in S\,\right)$. This is a general statistical tool, see Appendix D for a standalone formulation, discussion, and proof.
@@
$\mathtt{ERR}\left(\,\mathtt{mech}\mid\mathtt{adv}\,\right)$ $\geq\Omega\left(\,\frac{1}{|S|^{2}}\sum_{t\in S}\;\max\left\{\,\tfrac{1}{q_{t}},\;\sqrt{|S|}\,\right\}\,\right),\qquad\text{where }q_{t}:=\operatornamewithlimits{\mathbb{E}}\left[\,p_{t}(a_{0})\,\right]$
$\geq\Omega\left(\,\frac{1}{|S|^{2}}\;\left(\,\sqrt{|S|}\cdot|S_{\mathtt{L}}|+\sum_{t\in S_{\mathtt{H}}}\;\tfrac{1}{q_{t}}\,\right)\,\right),$ (3.7)
@@
By harmonic-arithmetic mean inequality,
$\sum_{t\in S_{\mathtt{H}}}\frac{1}{q_{t}}\geq|S_{\mathtt{H}}|^{2}/N_{\mathtt{H}}\geq|S_{\mathtt{H}}|^{2}\cdot\mathtt{bench}(\mathcal{P})/|S|.$
Plugging this back into (3.7) and noting that $|S_{\mathtt{L}}|+|S_{\mathtt{H}}|=|S|\geq T/2$, we have
$\mathtt{ERR}\left(\,\mathtt{mech}\mid\mathtt{adv}\,\right)$ $\geq\Omega\left(\,\frac{1}{|S|^{2}}\;\left(\,\sqrt{|S|}\cdot|S_{\mathtt{L}}|+|S_{\mathtt{H}}|^{2}\cdot\mathtt{bench}(\mathcal{P})/|S|\,\right)\,\right)\geq\Omega\left(\,\mathtt{bench}(\mathcal{P})/T\,\right).$
4 Heterogeneous agents
We now consider the general case of heterogeneous agents. Agent heterogeneity is two-fold: different agent types may have different outcome distributions for the same arm, and different subjective (and not directly observable) utilities for the same outcome. We model this as follows: public types determine outcome distributions and the private types determine subjective utilities. The intuition is that beliefs about objective outcomes are driven by the patient’s objective medical history, which is typically well-documented and available to the clinical trial. However, patients’ preferences over outcome distributions can be highly subjective, e.g., in evaluating the disutility of particular medical complications. These subjective preferences are not known to the mechanism. In our model, each agent reports its private type upon arrival, and the mechanism must incentivize truthful reporting (in addition to the BIR property).
4.1 Overview
We obtain matching upper and lower bounds like in Section 3, but with a new benchmark, mechanism, and analysis. This novelty is necessitated by the BIC condition which binds the types together. [^19] Indeed, without private types one can use mechanism from Section 3 tailored to each particular type. We define a new benchmark, denoted $\mathtt{bench}(\mathcal{P},\Theta)$, which generalizes (3.1) to a given set $\Theta$ of agent types. Compared to the expression in (3.1), this benchmark requires all BIC and BIR per-round policies under the $\inf$, and takes the $\max$ over $\Theta$ (as well as over all states). Our mechanism attains the same error bound as in Theorem 3.5 relative to $\mathtt{bench}(\mathcal{P},\Theta)$, with a similar but much more technical sufficient condition for BIR. The lower bound carries over from Theorem 3.7, replacing $\mathtt{bench}(\mathcal{P})$ with $\mathtt{bench}(\mathcal{P},\Theta)$.